OpenStudy (anonymous):

A scientist uses a scale Y for measuring temperature. In this scale, pure water melts at 10oY and boils at 130oY at standard atmospheric pressure. The scientist measures the temperature at which sodium melts to be 127oY. Express this temperature in degrees celcius

OpenStudy (anonymous):

130oY------> 100oC 10oY -------->0oc Therefore 1unit of oY=100/120=5/6 oC 127oY= 130oY - 3oY= 100oC- 3*5/6oY = (100-2.5)oY=97.5oC Final Ans 97.5oC

OpenStudy (anonymous):

@shivam_bhalla pls explain this line 127oY= 130oY - 3oY= 100oC- 3*5/6oY = (100-2.5)oY=97.5oC

OpenStudy (anonymous):

where u ve 3*5/6oY

OpenStudy (anonymous):

i got it now

OpenStudy (anonymous):

I am not sure whether this is right. Only applying logic. 127oY= 130oY - 3oY WKT 130oY = 100oC And WKT, 1oY=5/6oC Therefore, 3oY = 3 *(5/6)oC = 2.5 oC. Finally we get, 127oY= 100oC-2.5oC=97.5oC I still say that somebody verify this because I am not sure.

OpenStudy (anonymous):

it is right..